A normally distributed population has a mean of 76 and a standard deviation of 19. Determine the probability that a random sample of size 22 has an average between 72 and 76.

Round to four decimal places.
Answer #1

Solution :

Given that,

mean = \mu = 76

standard deviation = \sigma =19

n=22

\mu\bar x = 76

\sigma\bar x =\sigma  / \sqrt n = 19/ \sqrt 22=4.0508

= P(72< \bar x    < 76) = P[(72 - 76) / 4.0508< ( \bar x - \mu \bar x ) / \sigma \bar x < (76- 76) /4.0508 )]

= P( -0.99< Z <0 )

= P(Z <0 ) - P(Z <-0.99 )

Using z table,  

= 0.5 - 0.1611

= 0.3389